# Exposing Barriers to Flexibility Aggregation in Unbalanced Distribution Networks

### v2

@@ I Introduction
Modern distribution systems are evolving rapidly with the increasing integration of distributed energy resources (DER), particularly in low voltage (LV) networks [1, 2, 3]. Controllable DER, such as battery energy storage systems, prosumers, and electric vehicle aggregators, have the technical ability to adjust their power exchange with the grid and thus provide flexibility services, for example, by helping to balance electricity supply and demand [4, 5]. Such services can improve the operation of distribution systems and can also be aggregated and traded between distribution and transmission system operators (DSO and TSO) [6, 7, 8].
To enable flexibility services, flexibility aggregation models are essential both at the local level (within LV feeders) and system level (at primary substations or TSO-DSO interfaces) [9, 10, 11, 12, 13, 3, 14, 5, 15, 16, 17, 18, 19, 20]. These models aim to estimate the limits of flexible active (P) and reactive (Q) power exchanges at specific locations of a distribution network. By estimating the P-Q limits, DSO can clear flexibility markets while managing network constraints. The sets of all feasible power exchanges (operating points) in the P-Q space are known as aggregated P-Q flexibility areas or nodal operating envelopes [5, 14].
@@
  • Barrier #1: Lack of DER coordination. Most flexibility aggregation models assume that all flexibility providers (referred to as flexible units in this work) can be perfectly coordinated [9, 10, 11, 12, 5, 15, 16, 17, 18, 19, 20]. Such centralised aggregation methods imply that units at different locations can jointly adjust their power in order to manage voltage constraints and keep the network’s operation feasible. However, in real distribution systems, flexible units connected to different phases across the network may not be perfectly coordinated due to low observability, data exchange issues, single-phase connection, behind-the-meter devices, and other constraints [14, 13]. These practical limitations of DER coordination create a barrier to effective flexibility aggregation.
  • Barrier #2: Impacts of phase unbalance. Many existing studies rely on single-phase power flow models that implicitly assume perfectly balanced conditions in distribution networks [9, 10, 11, 12, 5, 14, 15, 16, 17, 18, 19, 20]. Although a few works on flexibility services do consider unbalanced distribution systems and apply three-phase optimal power flow (OPF) models, for example [2, 13, 3], they do not quantify the ensuing voltage unbalances and the volume of flexibility lost due to voltage unbalance limits.
@@
These barriers are present in real distribution systems as LV networks are inherently unbalanced and the increase of single-phase uncoordinated distributed generation will further aggravate existing voltage unbalances [21, 22]. For example, suboptimal selection of DER connection phases can lead to voltage unbalance and limit DER capacity [23, 24]. The need for simulating DER flexibility using accurate three-phase power flow models is acknowledged by the literature [25, 2, 26, 3, 27]. However, aside from the recent work [28], there have been no attempts to analyse the aggregated P-Q flexibility for unbalanced LV distribution networks. Note that although P-Q flexibility areas were calculated using a three-phase OPF model in [28], the impact of voltage unbalance constraints was not explicitly quantified. Moreover, no studies have analysed the loss of coordination between DER connected to different phases and its impacts on voltage unbalance.
@@
The models supporting the framework have been developed in Julia programming language and published as an open-source tool. The tool, named 3FlexAnalyser.jl, and the case studies are available at [29].
@@
The rest of the paper is organised as follows. Section II introduces the modelling framework for simulating DER operation in LV unbalanced networks and quantifying the impacts of voltage unbalance and phase coordination constraints. In Section III, extensive simulations are performed for an illustrative 5-bus distribution system and a real 221-bus LV network in the UK. Section IV further discusses the aspects of flexibility provision in unbalanced LV distribution networks. Section V concludes the paper.
II Modelling Framework
@@
To accurately simulate LV distribution networks with flexible units connected to different phases, a multi-phase flexibility estimation model (1a)-(1p) is formulated as a nonlinear programming problem (NLP). The formulation is based on the bus injection model widely applied in unbalanced distribution network optimisation studies [30, 31]. Variables of the model are defined for each phase $\phi$ and include complex voltages and complex power injections at each bus $i$, $V_{i,\phi}$, $s_{i,\phi}$, complex power flows for each line $(i,j)$, $s_{ij,\phi}$, active and reactive power of generators $P_{g,i,\phi}$, $Q_{g,i,\phi}$, and active and reactive power of flexible units $P_{f,i,\phi}$, $Q_{f,i,\phi}$. Note that the outputs of flexible units for each phase are the primary control variables of the optimisation problem used to construct the aggregated P-Q flexibility of the network. Voltage unbalances for the selected set of buses $\mathcal{N}^{\text{vu}}$ are explicitly modelled using the positive and negative voltage sequences, ${V_{1}}_{i},{V_{2}}_{i}$, and voltage unbalance factors (VUF), $VUF_{i}$. The original mathematical formulation of the three-phase power flow problem is adapted from the PowerModelsDistribution.jl package [31]. It is then extended by including flexible units, voltage unbalance limits, and phase coordination constraints.
The objective function (1a) minimises or maximises the network’s active and reactive power injections at the reference bus, where coefficients $\alpha^{p}_{i}$ and $\alpha^{q}_{i}$ are introduced to control the optimisation directions in the P-Q space. In this formulation, reference bus $i^{\text{ref}}$ is the location for which flexibility from multiple flexible units is aggregated. Thus, the objective function with adjustable coefficients $\alpha^{p}_{i}$, $\alpha^{q}_{i}$ allows to iteratively (point by point) estimate the limits of aggregated flexibility in the P-Q space at the reference bus [14, 32]. The role of the objective function (1a) in algorithms for constructing P-Q flexibility areas is further discussed later in this section, with a step-by-step illustration provided in Fig. 1. Equations (1b)-(1k) represent the nonlinear AC OPF constraints. Specifically, voltage limits are imposed in (1b), complex power flows and bus injections are defined in (1c), (1d), where $Y_{ij}$ is the admittance matrix. Active and reactive power balance constraints for each node are introduced in (1e), (1f), where $d\in\mathcal{D}$ represents loads of the network. Line capacity limits are included in (1g). Finally, active and reactive power of generators and flexible units is limited in (1h)-(1k).
@@
To characterise the flexibility services provided by LV distribution networks, the concept of P-Q flexibility areas is applied to the context of multi-phase flexibility provision. Mathematically, the operation limits of multiple flexible units can be described using the Minkowski addition of their P-Q capability sets [5, 37]. That is, the limits of the aggregated DER flexibility in the P-Q space provided at the reference bus can be estimated as the Minkowski addition of individual flexibility limits of all available flexible units. However, the Minkowski addition does not consider network constraints. To accurately estimate the aggregated DER flexibility considering network constraints, the formulated optimisation model (1a)-(1p) can be solved iteratively, and the boundary of the network flexibility area at the reference bus can be estimated with the desired level of granularity. Multiple algorithms have been proposed for constructing the boundaries of flexibility areas, including Monte Carlo simulations and various optimisation methods [38, 18, 19]. In this work, the $\epsilon$-constraint method is chosen as a simple and straightforward algorithm for approximating the P-Q flexibility area by considering $k$ intervals [16, 14, 32]. At each interval, the feasible P-Q space of the aggregated DER flexibility at the reference bus $i^{\text{ref}}$ is constrained by inequalities (2a) or (2b).
@@
Figure 1: Step-by-step illustration of the interval-based OPF algorithm used to construct the aggregated P-Q flexibility area for a single phase: (a) the initial operating point with no flexibility is calculated; (b) the bounds of the area are obtained by calculating Q minimum and maximum, points 1 and 2, and P minimum and maximum, points 3 and 4; (c) these limits are discretised into intervals – 10 Q intervals in this example; (d) the first Q interval is considered by constraining the Q component of the aggregated flexibility, the corresponding P limits (boundary points 5 and 6) are calculated; (e) after evaluating all intervals, the final area is estimated as the hull of all boundary points. The algorithm is executed separately for each phase.
@@
III Case Studies
In this section, the proposed framework is demonstrated using two case studies: (i) a simple 5-bus system with DER flexibility available at a single location; (ii) a real 221-bus distribution network in the UK with 12 flexible units. Data for these cases is available at [39, 29]. For each system, the limits of flexibility services are characterised via the concept of aggregated P-Q flexibility areas. Then, the impacts of voltage unbalance and phase coordination are quantified. Finally, a stochastic analysis is performed to evaluate how frequently voltage unbalance violations can occur due to uncoordinated flexible resources. All simulations have been performed with JuMP 1.11.0 for Julia 1.6.1 language and Ipopt 3.14.4 solver.
III-A Case Study: Illustrative 5-bus System

### v3

@@ I Introduction
Modern distribution systems are evolving rapidly with the increasing integration of distributed energy resources (DER), particularly in low voltage (LV) networks [Eid2016, Chen2020, Liu_Ochoa2022]. Controllable DER, such as battery energy storage systems (BESS), prosumers, and electric vehicles (EV), have the technical ability to adjust their power exchange with the grid and thus provide flexibility services, for example, by helping to balance electricity supply and demand [Schittekatte2020, Riaz2022]. Such services can improve the operation of distribution systems and can also be aggregated and traded between distribution and transmission system operators (DSO and TSO) [Gerard2018, Vicente-Pastor2019, Muller2019, Givisiez2020, Yi2021, Paredes2023].
To enable flexibility services, flexibility aggregation models are essential both at the local level (within LV feeders) and system level (at primary substations or TSO-DSO interfaces) [Heleno2015, Silva2018, Capitanescu2018, Contreras2018, Kalantar-Neyestanaki2020, Tan2020, Bolfek2021, Contreras2021, Lopez2021, Petrou2021, Sarstedt2022, Fruh2022, Nazir2022, Riaz2022, Liu_Ochoa2022, ChurkinSegmentation, Churkin2024, Rabiee2024, Chrysostomou2025, liu2025region, Pozo2025, Jiang2025]. These models aim to estimate the limits of flexible active (P) and reactive (Q) power exchanges at specific locations of a distribution network. By estimating the P-Q limits, DSO can clear flexibility markets while managing network constraints. The sets of all feasible power exchanges (operating points) in the P-Q space are known as aggregated P-Q flexibility areas or nodal operating envelopes [Riaz2022, Churkin2024].
@@
  • Barrier #1: Lack of DER coordination. Most flexibility aggregation models assume that all flexibility providers (referred to as flexible units in this work) can be perfectly coordinated [Yi2021, Heleno2015, Silva2018, Capitanescu2018, Contreras2018, Kalantar-Neyestanaki2020, Tan2020, Bolfek2021, Contreras2021, Lopez2021, Sarstedt2022, Fruh2022, Nazir2022, Riaz2022, ChurkinSegmentation, Rabiee2024, Chrysostomou2025, liu2025region, Jiang2025, Pozo2025]. Such centralised aggregation methods imply that units at different locations can jointly adjust their power in order to manage voltage constraints and keep the network’s operation feasible. However, in real distribution systems, flexible units connected to different phases across the network may not be perfectly coordinated due to low observability, data exchange issues, single-phase connection, behind-the-meter devices, and other constraints [Churkin2024, Petrou2021]. These practical limitations of DER coordination create a barrier to effective flexibility aggregation.
  • Barrier #2: Impacts of phase unbalance. Many existing studies rely on single-phase power flow models that implicitly assume perfectly balanced conditions in distribution networks [Yi2021, Heleno2015, Silva2018, Capitanescu2018, Contreras2018, Kalantar-Neyestanaki2020, Tan2020, Bolfek2021, Contreras2021, Lopez2021, Sarstedt2022, Fruh2022, Nazir2022, Riaz2022, ChurkinSegmentation, Churkin2024, Rabiee2024, Chrysostomou2025, liu2025region, Jiang2025, Pozo2025]. Although a few works on flexibility services do consider unbalanced distribution systems and apply three-phase optimal power flow (OPF) models, e.g. [Chen2020, LEIVA2020, FOBES2020, Petrou2021, Wang2021, Avramidis2021, Liu_Ochoa2022, Russell2023, Antic2024, Liu2024, Liu2024superellipsoid, Lu2024, Zhang2025], they do not quantify the ensuing voltage unbalances and the volume of flexibility lost due to voltage unbalance limits.
@@
These barriers are present in real distribution systems as LV networks are inherently unbalanced and the increase of single-phase uncoordinated distributed generation will further aggravate existing voltage unbalances [Ma2020, mousa2024comprehensive]. For example, suboptimal selection of DER connection phases can lead to voltage unbalance and limit DER capacity [Wang2021, Antic2024]. The need for simulating DER flexibility using accurate three-phase power flow models is acknowledged by the literature [LEIVA2020, Chen2020, Avramidis2021, Liu_Ochoa2022]. However, aside from the recent work [Bruno2021], there have been no attempts to analyse the aggregated P-Q flexibility for unbalanced LV distribution networks. Note that although P-Q flexibility areas were calculated using a three-phase OPF model in [Bruno2021], the impact of voltage unbalance constraints was not explicitly quantified. Moreover, no studies have analysed the loss of coordination between DER connected to different phases and its impacts on voltage unbalance. A detailed review of the existing literature and associated research gaps is provided in Section II.
@@
The models supporting the framework have been developed in Julia programming language and published as an open-source tool. The tool, named 3FlexAnalyser.jl, and the case studies are available at [3FlexAnalyser].
@@
The rest of the paper is organised as follows. Section II provides a literature review by mapping the relevant research directions and representative studies. Section III introduces the modelling framework for simulating DER operation in LV unbalanced networks and quantifying the impacts of voltage unbalance and phase coordination constraints. In Section IV, extensive simulations are performed for an illustrative 5-bus distribution system and a real 221-bus LV network in the UK. Section V further discusses the modelling aspects of flexibility aggregation. Section VI concludes the paper.
II Literature Review
The relevant research on flexibility in distribution networks can be broadly divided into two main directions. The first research stream focuses on aggregation models based on the concept of P-Q flexibility areas [Heleno2015, Silva2018, Capitanescu2018, Contreras2018, Kalantar-Neyestanaki2020, Tan2020, Bolfek2021, Contreras2021, Lopez2021, Sarstedt2022, Fruh2022, Riaz2022, ChurkinSegmentation, Churkin2024, Rabiee2024, Chrysostomou2025, liu2025region, Jiang2025, Pozo2025]. These areas characterise the feasible operating regions of distribution networks and enable the utilisation and trading of aggregated flexibility services between DSO and TSO. Most studies in this stream investigate the economic value of aggregated flexibility, exploring flexibility markets and coordination mechanisms. They rely on single-phase power flow models and therefore do not explicitly address flexibility provision in unbalanced LV networks. The second direction considers the operation of DER in unbalanced distribution networks, relying on three-phase OPF formulations to capture phase coupling and network constraints. However, these studies mainly focus on asset-level modelling, such as the analysis of DER operating envelopes or hosting capacity, and do not estimate aggregated P-Q flexibility areas at the network level. As a result, the integration of P-Q flexibility areas with three-phase modelling of unbalanced networks remains largely unexplored. This work is positioned at the intersection of the two research directions and addresses this gap. A detailed literature review mapping is shown in Fig. 1.
In the research direction on aggregation models for flexibility assessment, significant efforts have been devoted to the development of algorithms for computing the boundaries of flexibility areas in the P-Q space [Heleno2015, Silva2018, Capitanescu2018, Contreras2018, Bolfek2021, Contreras2021, Lopez2021]. A range of algorithms has been proposed, including random sampling and iterative optimisation-based methods. Several studies have extended P-Q flexibility aggregation to probabilistic settings by accounting for forecast uncertainty, stochastic DER dispatch, and the reliability of flexible units [Kalantar-Neyestanaki2020, Tan2020, ChurkinSegmentation]. Another group of works focuses on the economics of flexibility provision, developing models for cost-optimal aggregation and valuation of flexibility [Riaz2022, Sarstedt2022, Churkin2024, liu2025region]. Recent studies have further explored the properties of flexibility aggregation and disaggregation, proposing data-driven models and approximation approaches [Fruh2022, Rabiee2024, Chrysostomou2025, Jiang2025, Pozo2025]. However, research in this direction relies on single-phase power flow models and does not explicitly consider flexibility provision in unbalanced LV networks.
In parallel, multiple studies have investigated the flexible operation of DER in unbalanced distribution networks. This progress has been facilitated by advances in three-phase power flow modelling [Gan2014, Bernstein2018] and the release of open-source distribution power flow frameworks [FOBES2020]. Leveraging accurate three-phase power flow models, existing studies investigate the operation of DER in unbalanced distribution networks, primarily focusing on estimating network hosting capacity and DER operating limits [Chen2020, LEIVA2020, Petrou2021, Wang2021, Avramidis2021, Liu_Ochoa2022, Liu2024, Liu2024superellipsoid, Lu2024]. Several studies have examined the impact of phase unbalance on DER operation by explicitly accounting for voltage unbalance constraints [Russell2023, Antic2024, Zhang2025]. However, these works do not adopt the concept of aggregated P-Q flexibility areas and do not quantify the impact of phase coordination and voltage unbalance constraints on aggregated flexibility.
It follows that the intersection of these two research directions remains largely unexplored, as existing studies typically focus either on simplified single-phase modelling for P-Q flexibility aggregation or on detailed three-phase formulations without adopting the concept of aggregated P-Q flexibility areas. A notable emerging exception is [Bruno2021], which demonstrates that aggregated P-Q flexibility of an unbalanced network can be computed by accounting for flexible resources and network constraints in each phase. Yet, the impact of phase coordination and voltage unbalance constraints on three-phase P-Q flexibility areas is not explicitly quantified. This work further advances the modelling of aggregated P-Q flexibility areas and bridges this gap.
Figure 1: Literature review mapping. Columns represent the two research directions most relevant to this work. Representative studies are located according to their contributions to these directions and are arranged chronologically, with more recent publications displayed at the bottom. Each study is visualised by a circle with the corresponding reference number.
III Modelling Framework
@@
To accurately simulate LV distribution networks with flexible units connected to different phases, a multi-phase flexibility estimation model (1a)-(1p) is formulated as a nonlinear programming problem (NLP). The formulation is based on the bus injection model widely applied in unbalanced distribution network optimisation studies [Gan2014, FOBES2020, Bernstein2018]. Variables of the model are defined for each phase $\phi$ and include complex voltages and complex power injections at each bus $i$, $V_{i,\phi}$, $s_{i,\phi}$, complex power flows for each line $(i,j)$, $s_{ij,\phi}$, active and reactive power of generators $P_{g,i,\phi}$, $Q_{g,i,\phi}$, and active and reactive power of flexible units $P_{f,i,\phi}$, $Q_{f,i,\phi}$. Note that the outputs of flexible units for each phase are the primary control variables of the optimisation problem used to construct the aggregated P-Q flexibility of the network. Voltage unbalances for the selected set of buses $\mathcal{N}^{\text{vu}}$ are explicitly modelled using the positive and negative voltage sequences, ${V_{1}}_{i},{V_{2}}_{i}$, and voltage unbalance factors (VUF), $VUF_{i}$. The original mathematical formulation of the three-phase power flow problem is adapted from the PowerModelsDistribution.jl package [FOBES2020]. This three-phase AC OPF formulation inherently captures the propagation of unbalanced voltages and currents along network lines and allows the explicit modelling of single-phase, two-phase, and three-phase load connections. [^1] In this work, network data is prepared using the OpenDSS format with a three-phase three-wire AC OPF formulation. The neutral conductor and grounding are not explicitly modelled. Investigating four-wire network modelling is left for future research. It is then extended by including flexible units, voltage unbalance limits, and phase coordination constraints.
The objective function (1a) minimises or maximises the network’s active and reactive power injections at the reference bus, where coefficients $\alpha^{p}_{i}$ and $\alpha^{q}_{i}$ are introduced to control the optimisation directions in the P-Q space. In this formulation, reference bus $i^{\text{ref}}$ is the location for which flexibility from multiple flexible units is aggregated. Thus, the objective function with adjustable coefficients $\alpha^{p}_{i}$, $\alpha^{q}_{i}$ allows to iteratively (point by point) estimate the limits of aggregated flexibility in the P-Q space at the reference bus [Churkin2024, ChurkinSegmentation]. The role of the objective function (1a) in algorithms for constructing P-Q flexibility areas is further discussed later in this section, with a step-by-step illustration provided in Fig. 2. Equations (1b)-(1k) represent the nonlinear AC OPF constraints. Specifically, voltage limits are imposed in (1b), complex power flows and bus injections are defined in (1c), (1d), where $Y_{ij}$ is the admittance matrix. Active and reactive power balance constraints for each node are introduced in (1e), (1f), where $d\in\mathcal{D}$ represents loads of the network. Line capacity limits are included in (1g). Finally, active and reactive power of generators and flexible units is limited in (1h)-(1k).
@@
Figure 2: Conceptual step-by-step illustration of the interval-based OPF algorithm for constructing an aggregated P-Q flexibility area: (a) the initial operating point with no flexibility is calculated; (b) the bounds of the area are obtained by calculating Q minimum and maximum, points 1 and 2, and P minimum and maximum, points 3 and 4; (c) these limits are discretised into intervals – 10 Q intervals in this example; (d) the first Q interval is considered by constraining the Q component of the aggregated flexibility, the corresponding P limits (boundary points 5 and 6) are calculated; (e) after evaluating all intervals, the final area is estimated as the hull of all boundary points. The algorithm is executed separately for each phase. The P and Q axes in this figure are schematic and do not correspond to specific numerical values. In real calculations for distribution networks, these axes can be expressed in kW and kVAr or in per-unit values.
To characterise the flexibility services provided by LV distribution networks, the concept of P-Q flexibility areas is applied to the context of multi-phase flexibility provision. Mathematically, the operation limits of multiple flexible units can be described using the Minkowski addition of their P-Q capability sets [Riaz2022, AGARWAL2002, Churkin2024]. That is, the limits of the aggregated DER flexibility in the P-Q space provided at the reference bus can be estimated as the Minkowski addition of individual flexibility limits of all available flexible units. However, Minkowski addition assumes unconstrained DER operation and neglects any physical constraints. In practical applications, flexible units operate under network constraints, such as voltage and power flow limits, which makes direct Minkowski summation of individual P-Q capability sets invalid. Therefore, to accurately estimate the aggregated DER flexibility considering network constraints, the formulated optimisation model (1a)-(1p) can be solved iteratively, and the boundary of the network flexibility area at the reference bus can be estimated with the desired level of granularity. Multiple algorithms have been proposed for constructing the boundaries of flexibility areas, including Monte Carlo simulations and various optimisation methods [Lopez2021, Bolfek2021, Contreras2021]. In this work, the $\epsilon$-constraint method is chosen as a simple and straightforward algorithm for approximating the P-Q flexibility area by considering $k$ intervals [Capitanescu2018, Churkin2024, ChurkinSegmentation]. At each interval, the feasible P-Q space of the aggregated DER flexibility at the reference bus $i^{\text{ref}}$ is constrained by inequalities (2a) or (2b).
@@
Note that the proposed framework is formulated as a steady-state flexibility aggregation problem optimised at a given discrete time instance. Thus, it does not explicitly model intertemporal constraints and the dynamics of flexibility provision. Flexible resources with time and energy coupling, such as BESS and EV, would require additional state variables and constraints to represent their operation over a scheduling horizon, including state-of-charge dynamics and ramping limits. In a multi-period formulation, the aggregated P-Q flexibility area at a given time would depend on the selected horizon, initial conditions, as well as prior and future operating decisions. In this work, the steady-state framework is intentionally selected to isolate and clearly demonstrate the impact of network constraints, voltage unbalance, and phase coordination on aggregated flexibility. While extensions to multi-period formulations are conceptually straightforward, they are beyond the scope of this paper.
IV Case Studies
In this section, the proposed framework is demonstrated using two case studies: (i) a simple 5-bus system with DER flexibility available at a single location; (ii) a real 221-bus distribution network in the UK with 12 flexible units. For each system, the limits of flexibility services are characterised via the concept of aggregated P-Q flexibility areas. Then, the impacts of voltage unbalance and phase coordination are quantified. Finally, a stochastic analysis is performed to evaluate how frequently voltage unbalance violations can occur due to uncoordinated flexible resources.
Data for these case studies is publicly available at [churkin_Zenodo, 3FlexAnalyser] in the OpenDSS network file format. These files can be directly used in the developed 3FlexAnalyser.jl tool to reproduce the results presented in this section. The repository also provides a description of the input-output data structure, key modelling parameters, and the required Julia packages.
Note that in the considered case studies, thermal loading constraints of lines and transformers are not binding for the analysed operating points and flexibility ranges. The observed reductions in aggregated flexibility are therefore driven by voltage constraints rather than by line congestion. It is also assumed that the upstream network (supply source) is balanced at the point of connection. This assumption isolates the impact of downstream network constraints and voltage unbalance. Introducing upstream unbalance conditions would shift the feasible operating region, affecting the aggregated P-Q flexibility areas differently in each phase.
All simulations were performed using the JuMP modelling language in Julia and solved with the nonlinear optimisation solver IPOPT. This solver is widely used in the power systems literature for solving large-scale nonlinear OPF problems [Alizadeh2022]. For the considered case studies, IPOPT exhibited reliable convergence across the feasible operating points. While other solvers and flexibility aggregation algorithms can be used, they are not expected to alter the qualitative insights of this study.
IV-A Case Study: Illustrative 5-bus System